OpenStudy (erniellyd28):

kara's custom tees experienced fixed costs $200 and variable costs of $3 a shirt. writea n equation that can be used to determine the total expenses encountered by kara's custom tees. let x be the number of shirts, and let c(x) be the total cost of producing x shirts. then c(x)=? the cost of producing 8 shirts is $?

OpenStudy (mathstudent55):

cost of producing shirts = fixed costs + variable costs

OpenStudy (mathstudent55):

The fixed costs are $200. They are fixed costs because no matter how many shirts are made, the fixed costs remain at $200. The variable costs are variable because they depend on the number of shirts made. Each shirt has a cost of $3. How much would x shirts cost?

OpenStudy (erniellyd28):

$24?

OpenStudy (erniellyd28):

8 shirts times $3=24

OpenStudy (mathstudent55):

Yes, 8 shirt contribute a variable cost of $24 since 8 * $3 = $24 1 shirt has a variable cost of $3 * 1 = $3 5 shirts have a variable cost of $3 * 5 = $15 The unknown number of shirts, x, has what variable cost?

OpenStudy (erniellyd28):

200+3x

OpenStudy (mathstudent55):

Correct. C(x) = 200 + 3x Now to find the cost of making 8 shirts, plug in 8 for x in the function above.

OpenStudy (erniellyd28):

224?

OpenStudy (mathstudent55):

Correct.

OpenStudy (erniellyd28):

i'm smart lol

OpenStudy (erniellyd28):

jk

OpenStudy (mathstudent55):

I see. Good job!

OpenStudy (erniellyd28):

thank you very much for your help :-)

OpenStudy (mathstudent55):

You're welcome.